OpenStudy (samara8954):

Which equation could be used to solve the problem? In the summer, Ynez earns $15 per week mowing lawns. At that rate, if she works 12 weeks, how much will she make (m)?

OpenStudy (samara8954):

m = 12 ÷ 15 B. m = 15 + 12 C. m = 15 ÷ 12 D. m = 15 • 12

OpenStudy (samara8954):

÷ 15 B. m = 15 + 12 C. m = 15 ÷ 12 D. m = 15 • 12

OpenStudy (michele_laino):

please if Ynez works only one week, how does she earn?

OpenStudy (michele_laino):

@Samara8954

OpenStudy (samara8954):

I don"t know

OpenStudy (michele_laino):

please read text of your problem

OpenStudy (michele_laino):

please retry!

OpenStudy (samara8954):

add

OpenStudy (michele_laino):

no, please you have to multiply, so D.

OpenStudy (michele_laino):

if Ynez earns 15 $ for week, she earns 30$ every two weeks, and...she earns 15*12=180$ after 12 weeks!

OpenStudy (michele_laino):

is it all clear?

OpenStudy (michele_laino):

please?